OpenStudy (anonymous):

Factor the four-term polynomial by grouping. Another help please? 3q^2+8pq-3q-8p

OpenStudy (anonymous):

q(3q+8p)-(3q+8p) =(q-1)(3q+8p)

OpenStudy (anonymous):

3q(q-1) + 8p(q-1) = (3q+8)(q-1)

OpenStudy (anonymous):

thank you very much guys

OpenStudy (anonymous):

yw